The sum of the squares of two consecutive integers is 85. What are the integers?
1
Expert's answer
2019-07-30T13:10:16-0400

Let these integers be n and (n+1). We have:


"n^2+(n+1)^2=85"

"n^2+n^2+2n+1=85"

"2n^2+2n-84=0"

"n^2+n-42=0"

We need to solve this quadratic equation:


"D=1^2-4(1)(-42)=169=13^2"

"n_1=\frac{-1-13}{2(1)}=-7"

"n_2=\frac{-1+13}{2(1)}=6"

So, we hawe two pairs of consecutive integers:


"(-7,-6)\ and \ (6,7)"